OpenStudy (anonymous):

A recipe that makes 7 servings calls for 1/2 cup of apple juice and 2/3 cup of cranberry juice. How many cups of juice are there per serving? A. 1/6 B. 3/5 C. 1 1/16 D. 8 1/6

OpenStudy (texaschic101):

apple juice + cranberry juice = total juice 1/2 + 2/3 = total juice 3/6 + 4/6 = total juice 7/6 = total juice so 7 servings calls for 7/6 cups of juice..... 7/(7/6) = 1/x --- (7 serv to 7/6 cup of j = 1 serv to x cups of j) cross multiply (7)(x) = (7/6)(1) 7x = 7/6 x = (7/6) / 7 x = 7/6 * 1/7 x = 7/42 which reduces to 1/6 so 1 servings calls for 1/6 cups of juice any questions ?

OpenStudy (anonymous):

thanks @texaschic101